# A binary neutron star merger search pipeline powered by deep learning

### v1

@@ V.1 Injection run results
We measure our pipeline’s performance using an injection run, which is a search for simulated signals placed into noise. Our injection run data is the set of BNS injections from the GWTC-3 catalogue that was used to estimate the offline detection pipeline sensitivities in O3 [5, 57]. The parameters of these injections can be found in Table 3. Since we used the third week of O3 for collecting our background, we selected the fourth week of O3 for our injection run. After applying the same cuts to the data as described in Sec. II.3, we were left with a set of 2,800 injections to test our search’s sensitivity with. The injection run was performed using the same procedure described in Sec. IV, with the only changes being adding the injections to the noise before matched filtering, and the lack of time shifts. We then assign a false alarm rate to the injections using the collected background, or the Gaussian extrapolation if the ranking statistic is greater than the highest background point.
@@
Polarisation Uniform [0, $\pi$]
Network SNR - $\geq$ 6
@@
The next step is to determine when a BNS merger has been detected by the pipeline. We consider an injection detected, or ‘found’, if there is a trigger within one second of the injection time below a false alarm rate threshold of 1 per 2 months ($\sim 2\times 10^{-7}$ Hz), the O3 open public alert threshold [5]. Figure 4 shows the distribution of events that were missed or found at this threshold, with respect to their effective distance in the most sensitive detector, Livingston, and detector-frame chirp mass. As expected, our pipeline tends to find closer events, and is able to find events with a higher chirp mass out to further distances, since heavier binaries have a larger GW amplitude than lighter binaries at the same distance. The pipeline finds 50% of all events closer than an effective distance of 530 Mpc, and 100% of all events closer than an effective distance of 175 Mpc. The closest missed event had a recovered Livingston SNR that was 50% higher than its injected SNR due to a loud blip glitch. This indicates there is a limit to how loud a glitch can be before the pipeline is unable to recover the event.
Figure 4: Left: Missed and found injections against their detector-frame chirp mass and effective distance in the Livingston detector. An injection is classed as ‘found’ if its assigned false alarm rate is less than 1 per 2 months. Right: Fraction of injections recovered closer than a given Livingston effective distance.
Figure 5 shows the recovered Hanford and Livingston SNR of the detected events. The recovered SNR for each detector is close to the injected SNR, as expected, where the injected SNR is the optimal matched filtering SNR of a signal with the given injection parameters. The spread at low SNRs is due to noise power having a greater influence on the recovered SNR, so individual noise realisations can cause the recovered SNR to noticeably increase or decrease. Six outlier events had a significantly higher recovered Hanford SNR than their injected SNRs. In all of these events, there was a loud blip glitch in the Hanford detector between 1 and 25 seconds before the merger time. This indicates that the neural network is capable of making detections even when a glitch significantly affects the SNR time series. However, the glitchy event missed at $\sim$ 175 Mpc shows that it is possible for loud signals to be missed due to loud glitches.
@@ V.1.1 Search sensitivity
where $N_{\text{draw}}$ is the number of drawn injections (rejected injections count towards $N_{\text{draw}}$), $N_{\text{det}}$ is the number of injections with $\mathcal{F}_{i}\leq\mathcal{F}$, $p_{\text{draw}}$ is the draw probability of an injection, and $p(\theta_{i})$ is the parameter distribution used [57]. The Monte Carlo uncertainty of the sensitive volume estimate is given by [60]
@@
To compare our method to the detection pipelines, we take the pipeline’s reported false alarm rates for the set of 2,800 injections and calculate their sensitive distances using Eqn. 6. We only compare to the PyCBC, GstLAL and MBTA pipelines, as the SPIIR pipeline did not run offline in O3, and the unmodelled cWB pipeline is not sensitive to BNS mergers. As we do not consider periods where one or both of the LIGO detectors are offline, the following sensitivities can be interpreted as the sensitive distance of the pipelines and our method when the Hanford and Livingston detectors are observing. One caveat for the pipeline sensitivities is that in O3 the existing pipelines used an additional metric, $p_{\text{astro}}$, to determine the likelihood that a trigger is of astrophysical origin. The sensitivity results reported in [5] only focus on triggers with $p_{\text{astro}}>0.5$, but since we do not compute a $p_{\text{astro}}$, we consider the false alarm rates for all of the pipelines irrespective of their estimated $p_{\text{astro}}$ values. Another caveat is that the other search pipelines also search for BBH and NSBH events, which slightly reduces their sensitivity to BNS events (further discussed in Sec. V.1.2). Figure 7 shows the two-detector sensitive distances of the offline detection pipelines and our method’s sensitive distance during the week 4 injection run. We find that our pipeline has a similar sensitivity to the PyCBC and MBTA pipelines above the 1 per 2 months detection threshold, and a similar sensitivity to the PyCBC and GstLAL pipelines below the detection threshold.
@@
While the sensitive distance of our method is comparable to the online pipelines’ sensitivities at all false alarm rates, this metric does not completely explore the expected BNS sensitivity improvement of adding our pipeline to the set of detection pipelines. Since in a real search an event is considered detected when found by one or more pipelines, and the pipelines do not necessarily detect the same events as each other, a useful quantity is the expected sensitivity increase from adding our pipeline to the set of detection pipelines, which is shown in Fig. 8. We find a 12% increase in the total number of events detected at the 1 per 2 months detection threshold from adding signals detected by only our pipeline to those detected by the three other pipelines. This increase in sensitivity comes from 124 events that were only detected by our pipeline. This compares favourably to the other pipelines, with PyCBC contributing the most single-pipeline detections (78) before considering our pipeline’s detections. Additionally, our pipeline detects several events that were previously only detected by one pipeline, and so increases the number of joint detections by 7%.
@@ V.1.2 Comparison to a PyCBC BNS-only search
A caveat for this sensitivity comparison is that the offline pipelines search for BNS, NSBH and BBH events, while our pipeline is currently only searching for BNS events. Searching over a larger parameter space allows for the detection of more sources, but reduces a search’s sensitivity to individual source types due to the greater rate of false alarms. Were our pipeline to also search for NSBH and BBH mergers, its BNS sensitivity would be slightly lower.
@@ V.2.1 GW170817
GW170817 is an important test of our pipeline for two reasons. Firstly, since GW170817 is in O2 but our neural network was trained on O3 data, it allows us to test whether our pipeline is capable of making detections despite the changes in detector PSD between observing runs. Secondly, GW170817 is notable for having a loud blip glitch present $\sim$1 second before the merger in the Livingston detector. Searching for GW170817 without removing the blip glitch beforehand will demonstrate whether our pipeline is capable of making real detections without glitch mitigation.
To test our pipeline’s ability to detect GW170817, we first performed a 1 week background run in O2, ending 512 seconds before the merger time of GW170817. We then performed a 1 week search immediately after the background run, such that GW170817 was present 512 seconds into the search. Triggers from this search were then assigned false alarm rates with the O2 background run, which are shown in Fig. 10. With this O2 background, GW170817 was assigned a false alarm rate of $1.1\times 10^{-12}$ Hz, or 1 per 30,000 years. The triggering template recovered a network SNR of 29.6 and has a detector frame chirp mass of 1.1978 M, which is consistent with the previously reported detector frame chirp mass of $1.1977^{+0.0008}_{-0.0003}$ M[6]. The recovered SNR and assigned false alarm rate are also similar to those of the detection pipelines, as shown in Table 4. We also find that the neural network’s predictions are insensitive to the glitch present in the Livingston detector, as shown in Fig. 11. The neural network’s predictions increase when GW170817 enters the input window, but do not change when the glitch enters the input window. From this, we find that our pipeline is confidently detecting GW170817, and is able to detect it despite the loud glitch.
@@
This work 29.6 $1.1\times 10^{-12}$
@@ V.2.2 GW190425
While GW170817 is relatively straightforward to detect with our method despite the glitch, GW190425 is more challenging. GW190425 was detected as a single detector event by the GstLAL pipeline, as the Hanford detector was offline and the Virgo SNR was below the detection threshold [7]. Our neural network is only trained on two detector events, and while the Virgo detector was online, we trained the neural network specifically on the two LIGO detectors. However, since we use an addition layer to merge the H and L branches of the neural network (see Fig. 1), by setting the output of the H branch to zeroes, the only contribution to the combiner sub-network will be from the L branch, thus approximating a single-detector pipeline. While we did not train our neural network on the single-detector case, and analysing events in this way was not considered when constructing our method, it provides a compelling proof-of-concept for further investigation into a single-detector extension of our pipeline.
@@ V.3 Sensitivity over O3
During online and offline searches, the CBC detection pipelines update their noise background over time to compensate for the non-stationarity of the interferometer noise. Since our neural network was trained on the first week of O3 and the background was collected in the third week of O3, we evaluate if there is any noticeable sensitivity loss throughout O3 to determine if the background needs updating or if the network needs retraining on long timescales. Since it would be difficult to determine if any changes in the pipeline’s sensitivity week-to-week are due to changes in interferometer sensitivity or due to the neural network performing poorly on noise it was not trained on, we compare the pipeline to the other offline pipelines’ sensitivities over O3. For this comparison, we perform additional week-long injection runs using the O3 injection set at a cadence of roughly one per month. Figure 14 shows the sensitive distance of the pipelines over the O3 dataset. Any sensitivity changes in our pipeline are also evident in the other detection pipelines, showing that our pipeline’s sensitivity is largely unchanged over the course of O3.
Figure 14: Sensitive distance of our pipeline compared to the offline detection pipelines at the detection FAR threshold of 1 per 2 months over the course of O3. The dark grey region is the break between O3a and O3b, during which the detectors were offline. The light grey regions are the weeks in which the injection runs were performed with our pipeline.
@@ VI Conclusions
Motivated by the results of this study, we aim to apply this method to NSBH mergers in the future. Detecting NSBH mergers with this method would not require any significant changes, apart from the template bank generation method used, as geometric banks are not capable of efficiently covering higher mass regions. A combined model capable of detecting BBH [22], BNS and NSBH mergers could then be investigated, as this would bring our detection method more in line with the current CBC pipelines which detect all three classes of events, and would improve the accuracy of our comparison to the existing pipelines. We will also investigate the pre-merger detection of BNS and NSBH mergers with this method, with the aim of contributing to early warning triggers to aid multi-messenger astronomy with gravitational waves.
@@ VII Data and Software Availability
The code we developed for our sample generation is available at https://github.com/alistair-mcleod/GWSamplegen. The code we developed for collecting our background and running our search is available at https://github.com/alistair-mcleod/infernus. These repositories can be used to reproduce the results presented in this work.
VIII Acknowledgements
@@
This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.

### v2

@@ V.1 Injection run results
We measure our pipeline’s performance using an injection run, which is a search for simulated signals placed into noise. Our injection run data is the set of BNS injections from the GWTC-3 catalog that was used to estimate the offline detection pipeline sensitivities in O3 [5, 57]. The parameters of these injections can be found in Table 3. Since we used the third week of O3 for collecting our background, we selected the fourth week of O3 for our injection run. After applying the same cuts to the data as described in Sec. II.3, we were left with a set of 2,800 injections to test our search’s sensitivity with. The injection run was performed using the same procedure described in Sec. IV, with the only changes being adding the injections to the noise before matched filtering, and the lack of time shifts. We then assign a false alarm rate to the injections using the collected background, or the Gaussian extrapolation if the ranking statistic is greater than the highest background point.
@@
Polarization Uniform [0, $\pi$]
Network SNR $\cdots$ $\geq$ 6
@@
The next step is to determine when a BNS merger has been detected by the pipeline. We consider an injection detected, or “found”, if there is a trigger within one second of the injection time below a false alarm rate threshold of 1 per 2 months ($\sim 2\times 10^{-7}$ Hz), the O3 open public alert threshold [5]. Figure 4 shows the distribution of events that were missed or found at this threshold, with respect to their effective distance in the most sensitive detector, Livingston, and detector-frame chirp mass. As expected, our pipeline tends to find closer events, and is able to find events with a higher chirp mass out to further distances, since heavier binaries have a larger GW amplitude than lighter binaries at the same distance. As shown in Fig. 4, the pipeline finds 34% of all injections in the dataset, 50% of all events closer than an effective distance of 530 Mpc, and 100% of all events closer than an effective distance of 175 Mpc. The closest missed event had a recovered Livingston SNR that was 50% higher than its injected SNR due to a loud blip glitch. This indicates there is a limit to how loud a glitch can be before the pipeline is unable to recover the event.
Figure 4: Left: missed and found injections against their detector-frame chirp mass and effective distance in the Livingston detector. An injection is classed as found if its assigned false alarm rate is less than 1 per 2 months. Right: fraction of injections that were found at or below a given Livingston effective distance.
Figure 5 shows the recovered Hanford and Livingston SNR of the detected events. The recovered SNR for each detector is close to the injected SNR, as expected, where the injected SNR is the optimal matched filtering SNR of a signal with the given injection parameters. The spread at low SNRs is due to noise power having a greater influence on the recovered SNR, so individual noise realizations can cause the recovered SNR to noticeably increase or decrease. Six outlier events had a significantly higher recovered Hanford SNR than their injected SNRs. In all of these events, there was a loud blip glitch in the Hanford detector between 1 and 25 seconds before the merger time. This indicates that the neural network is capable of making detections even when a glitch significantly affects the SNR time series. However, the glitchy event missed at $\sim$ 175 Mpc shows that it is possible for loud signals to be missed due to loud glitches.
@@ V.1.1 Search sensitivity
where $N_{\text{draw}}$ is the number of drawn injections (rejected injections count toward $N_{\text{draw}}$), $N_{\text{det}}$ is the number of injections with $\mathcal{F}_{i}\leq\mathcal{F}$, $p_{\text{draw}}$ is the draw probability of an injection, and $p(\theta_{i})$ is the parameter distribution used [57]. The Monte Carlo uncertainty of the sensitive volume estimate is given by [60]
@@
To compare our method to the detection pipelines, we take the pipeline’s reported false alarm rates for the set of 2,800 injections and calculate their sensitive distances using Eq. (6). We only compare to the PyCBC, GstLAL and MBTA pipelines, as the SPIIR pipeline did not run offline in O3, and the unmodeled cWB pipeline is not sensitive to BNS mergers. As we do not consider periods where one or both of the LIGO detectors are offline, the following sensitivities can be interpreted as the sensitive distance of the pipelines and our method when the Hanford and Livingston detectors are observing. One caveat for the pipeline sensitivities is that in O3 the existing pipelines used an additional metric, $p_{\text{astro}}$, to determine the likelihood that a trigger is of astrophysical origin. The sensitivity results reported in [5] only focus on triggers with $p_{\text{astro}}>0.5$, but since we do not compute a $p_{\text{astro}}$, we consider the false alarm rates for all of the pipelines irrespective of their estimated $p_{\text{astro}}$ values. Another caveat is that the other search pipelines also search for BBH and NSBH events, which slightly reduces their sensitivity to BNS events (further discussed in Sec. V.1.2). Figure 7 shows the two-detector sensitive distances of the offline detection pipelines and our method’s sensitive distance during the week 4 injection run. We find that our pipeline has a similar sensitivity to the PyCBC and MBTA pipelines above the 1 per 2 months detection threshold, and a similar sensitivity to the PyCBC and GstLAL pipelines below the detection threshold.
@@
While the sensitive distance of our method is comparable to the online pipelines’ sensitivities at all false alarm rates, this metric does not completely explore the expected BNS sensitivity improvement of adding our pipeline to the set of detection pipelines. Since in a real search an event is considered detected when found by one or more pipelines, and the pipelines do not necessarily detect the same events as each other, a useful quantity is the expected sensitivity increase from adding our pipeline to the set of detection pipelines, which is shown in Fig. 8. We find a 12% increase in the total number of events detected at the 1 per 2 months detection threshold from adding signals detected by only our pipeline to those detected by the three other pipelines. This increase in sensitivity comes from 124 events that were only detected by our pipeline. This compares favorably to the other pipelines, with PyCBC contributing the most single-pipeline detections (78) before considering our pipeline’s detections. Additionally, our pipeline detects several events that were previously only detected by one pipeline, and so increases the number of joint detections by 7%.
@@ V.1.2 Comparison to a PyCBC BNS-only search
A caveat for this sensitivity comparison is that the offline pipelines search for BNS, NSBH, and BBH events, while our pipeline is currently only searching for BNS events. Searching over a larger parameter space allows for the detection of more sources, but reduces a search’s sensitivity to individual source types due to the greater rate of false alarms. Were our pipeline to also search for NSBH and BBH mergers, its BNS sensitivity would be slightly lower.
@@ V.2.1 GW170817
GW170817 is an important test of our pipeline for two reasons. First, since GW170817 is in O2 but our neural network was trained on O3 data, it allows us to test whether our pipeline is capable of making detections despite the changes in detector PSD between observing runs. Second, GW170817 is notable for having a loud blip glitch present $\sim$1 second before the merger in the Livingston detector. Searching for GW170817 without removing the blip glitch beforehand will demonstrate whether our pipeline is capable of making real detections without glitch mitigation.
To test our pipeline’s ability to detect GW170817, we first performed a 1 week background run in O2, ending 512 seconds before the merger time of GW170817. We then performed a 1 week search immediately after the background run, such that GW170817 was present 512 seconds into the search. Triggers from this search were then assigned false alarm rates with the O2 background run, which are shown in Fig. 10. With this O2 background, GW170817 was assigned a false alarm rate of $6.9\times 10^{-13}$ Hz, or 1 per 46,000 years. Using the method for estimating the uncertainty of our FAR extrapolation described in Sec. IV, we find that GW170817’s FAR is between $1.7\times 10^{-12}$ Hz and $2.4\times 10^{-14}$ Hz. GW170817 would be considered confidently detected at any FAR in this range. The triggering template recovered a network SNR of 31.3 and has a detector frame chirp mass of 1.1978 M, which is consistent with the previously reported detector frame chirp mass of $1.1977^{+0.0008}_{-0.0003}$ M[6]. The recovered SNR and assigned false alarm rate are also similar to those of the detection pipelines, as shown in Table 4. We also find that the neural network’s predictions are insensitive to the glitch present in the Livingston detector, as shown in Fig. 11. The neural network’s predictions increase when GW170817 enters the input window, but do not change when the glitch enters the input window. From this, we find that our pipeline is confidently detecting GW170817, and is able to detect it despite the loud glitch.
@@
This work 31.3 $6.9\times 10^{-13}$
@@ V.2.2 GW190425
While GW170817 is relatively straightforward to detect with our method despite the glitch, GW190425 is more challenging. GW190425 was detected as a single detector event by the GstLAL pipeline, as the Hanford detector was offline and the Virgo SNR was below the detection threshold [7]. Our neural network is only trained on two detector events, and while the Virgo detector was online, we trained the neural network specifically on the two LIGO detectors. However, since we use an addition layer to merge the H and L branches of the neural network (see Fig. 1), by setting the output of the H branch to zeroes, the only contribution to the combiner subnetwork will be from the L branch, thus approximating a single-detector pipeline. While we did not train our neural network on the single-detector case, and analysing events in this way was not considered when constructing our method, it provides a compelling proof-of-concept for further investigation into a single-detector extension of our pipeline.
@@ V.3 Sensitivity over O3
During online and offline searches, the CBC detection pipelines update their noise background over time to compensate for the nonstationarity of the interferometer noise. Since our neural network was trained on the first week of O3 and the background was collected in the third week of O3, we evaluate if there is any noticeable sensitivity loss throughout O3 to determine if the background needs updating or if the network needs retraining on long timescales. Since it would be difficult to determine if any changes in the pipeline’s sensitivity week-to-week are due to changes in interferometer sensitivity or due to the neural network performing poorly on noise it was not trained on, we compare the pipeline to the other offline pipelines’ sensitivities over O3. For this comparison, we perform additional week-long injection runs using the O3 injection set at a cadence of roughly one per month. Figure 14 shows the sensitive distance of the pipelines over the O3 dataset. Any sensitivity changes in our pipeline are also evident in the other detection pipelines, showing that our pipeline’s sensitivity is largely unchanged over the course of O3.
Figure 14: Sensitive distance of our pipeline compared to the offline detection pipelines at the detection FAR threshold of 1 per 2 months over the course of O3. The dark gray region is the break between O3a and O3b, during which the detectors were offline. The light gray regions are the weeks in which the injection runs were performed with our pipeline.
@@ VI Conclusions
Motivated by the results of this study, we aim to apply this method to NSBH mergers in the future. Detecting NSBH mergers with this method would not require any significant changes, apart from the template bank generation method used, as geometric banks are not capable of efficiently covering higher mass regions. A combined model capable of detecting BBH [22], BNS and NSBH mergers could then be investigated, as this would bring our detection method more in line with the current CBC pipelines which detect all three classes of events, and would improve the accuracy of our comparison to the existing pipelines. We will also investigate the pre-merger detection of BNS and NSBH mergers with this method, with the aim of contributing to early warning triggers to aid multimessenger astronomy with gravitational waves.
@@ VII Data and Software Availability
The code we developed for our sample generation is available at [61]. The code we developed for collecting our background and running our search is available at [62]. These repositories can be used to reproduce the results presented in this work.
VIII Acknowledgments
@@
This research has made use of data or software obtained from the Gravitational Wave Open Science Center [63], a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.